OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

i am completely lost

OpenStudy (loser66):

\[cot x sec^4 x = cot x +2tan x +tan^3x\\or~ cotx sec (4x)= cotx +2 tan x + tan (3x)\]

OpenStudy (anonymous):

first one

OpenStudy (loser66):

just sec^4 = (1+tan^2)^2 and open it, then time cot into the expanding, and.... got the RHS

OpenStudy (anonymous):

what?

OpenStudy (psymon):

Yeah, once you make it into (1+tan^2(x))^2, you'll be switching your cot and your tan into sines and cosines to solve from there.

OpenStudy (jdoe0001):

@nickolgee429 have you covered yet binomial theorem?

OpenStudy (anonymous):

is that just separating the binomial into two parts? @jdoe0001

OpenStudy (jdoe0001):

into 3 parts

OpenStudy (jdoe0001):

a perfect square \(\bf \text{trinomial}\)

OpenStudy (loser66):

@Psymon no need to turn to sin/cos

OpenStudy (anonymous):

but its to the power of 2? so wouldn't i make (1+tan^2(x))^2 into (1+tan^2(x))x(1+tan^2(x))

OpenStudy (psymon):

Probably not, but it makes it visually easier to see if you're not used to it. I'm obviously no expert at it, but it lets me know for sure what I'm doing. It's a safe bet for me personally.

OpenStudy (jdoe0001):

@nickolgee429 yes it does, as @Loser66 correctly pointed out, expand that, and then multiply by cot(x)

OpenStudy (jdoe0001):

hmmm one sec.... it doesn't expand in to THAT

OpenStudy (loser66):

ok, let me show my work.

OpenStudy (jdoe0001):

@nickolgee429 I'm assuming you know what a perfect square trinomial is?

OpenStudy (anonymous):

like this: cot (x) times (1+tan^2(x)) times(1+tan^2(x))? and yes i know what a perfect square trinomial is.

OpenStudy (loser66):

\[cot (1 + tan^2)^2= cot (1 + 2 tan^2 + tan^4) = cot +\frac {2tan^2}{tan}+ \frac {tan^4}{tan}\] cot +2 tan+ tan^3

OpenStudy (jdoe0001):

hmmm http://statisticslectures.com/images/trinomial11.gif

OpenStudy (loser66):

because cot = \(\dfrac {1}{tan}\)

OpenStudy (anonymous):

ohh cot = \[\frac{ 1 }{ \tan }\]

OpenStudy (anonymous):

ok now i get it! Thanks everyone!

OpenStudy (loser66):

yw